# Equivariant Neural Networks for General Linear Symmetries on Lie Algebras

### v1

@@ F.1 Group Action Equivariance via Embedding Map
Theorem F.1 (Adjoint Equivariance).
The embedding $\varphi$ correctly models the Lorentz group action. For any $p\in\mathbb{R}^{4}$ and Lorentz transformation $\Lambda\in\operatorname{SO}(1,3)$, let $G=\operatorname{diag}(\Lambda,1)\in\operatorname{GL}(5)$. The map is equivariant in the sense that the standard action on $p$ corresponds to the adjoint action on $\varphi(p)$:
@@
Remark 3 (Generalization to Orthogonal Groups).
This embeddinging technique is not limited to the Lorentz group and can be readily generalized to any orthogonal group $\operatorname{O}(n)$ or special orthogonal group $\operatorname{SO}(n)$. For instance, in applications involving 3D point clouds where the symmetry is $\operatorname{SO}(3)$, a vector $p\in\mathbb{R}^{3}$ would be embedded into the Lie algebra $\mathfrak{gl}(4)$ as:
@@
F.2 Experimental Implementation
Dataset
@@ F.2 Experimental Implementation
Model
Our model leverages the established architecture of LorentzNet  (Gong et al., 2022), utilizing its stack of Lorentz Group Equivariant Blocks (LGEBs) for message passing on the jet’s particle cloud. While the original LorentzNet computes these features directly from the 4-momenta using the Minkowski inner product, our approach introduces a modified bilinear form based feature extraction. We first embed each pair of 4-momenta, $p_{i}$ and $p_{j}$, from the Minkowski space $\mathbb{R}^{1,3}$ into the Lie algebra $\mathfrak{gl}(5)$ via the map $p\mapsto X(p)$. The invariant features for the message passing are then derived from the bilinear form, $B(\cdot,\cdot)$, on this Lie algebraic space. The edge message $m_{ij}$ is thus constructed as:
$m_{ij}=\phi_{e}\Big(h_{i},h_{j},\psi\big(\widetilde{B}(X(p_{i}),X(p_{i}))\big),\psi\big(\widetilde{B}(X(p_{i}),X(p_{j}))\big)\Big)$ (38)
@@
Training Setup
For a fair comparison, our training procedure closely follows the protocol established in the LorentzNet (Gong et al., 2022). The model was trained for a total of 35 epochs on a single NVIDIA RTX 4090 GPU. We used the AdamW optimizer with a weight decay of 0.01 and a batch size of 128, matching the total effective batch size from the reference work. The learning rate was managed by the paper’s specific three-stage schedule: a 4-epoch linear warm-up to an initial rate of $1\times 10^{-3}$, followed by a 28-epoch CosineAnnealingWarmRestarts schedule, and a final 3-epoch exponential decay. After each epoch, the model with the highest validation accuracy was saved for final evaluation on the test set.
Appendix G Drone Experiment Details
@@
G.1 Geometric Framework for Equivariant Covariance Processing
@@
Theorem G.1 (Equivariance of the Logarithmic Map).
@@
G.3 Baseline and Model Implementation Details
@@
Eigendecomposition-based $\mathrm{SO}(3)$-Equivariant Baseline.
@@
The ReLN model shares the same base architecture as the VN model but incorporates ReLN-Bracket layer as an additional source of nonlinearity after the initial feature extraction block. In contrast to the VN backbone, ReLNs provide a unified framework for velocity and covariance processing. Velocities $\mathbf{v}\in\mathbb{R}^{3}$ are lifted into the Lie algebra $\mathfrak{so}(3)$ via $K=\mathbf{v}^{\wedge}$, while covariance $C$ (or $\log C$) is treated as a structured geometric input. Both transform under the same adjoint action: $K^{\prime}=RKR^{\top}$ and $C^{\prime}=RCR^{\top}$, enabling joint equivariant processing. Although $\mathrm{SPD}(n)$ is not a Lie algebra or group, it is a subset of $\mathrm{GL}(n)$. By taking the matrix logarithm, $\log C\in\mathrm{Sym}(n)\subset\mathfrak{gl}(n)$, covariances are embedded into a linear subspace compatible with Lie-algebra processing.
The network $\Phi$ fuses these inputs into a single matrix $A\in\mathbb{R}^{3\times 3}$, from which we extract the velocity estimate:
$A_{\mathrm{skew}}\coloneqq\tfrac{1}{2}(A-A^{\top})\in\mathfrak{so}(3),\qquad\hat{\mathbf{v}}=\text{Vee}(A_{\mathrm{skew}}).$ (46)
The extracted velocity is provably $\mathrm{SO}(3)$-equivariant by construction. See Appendix H for the full statement and proof.
G.4 Training and Evaluation Protocol
Problem Formulation
@@ G.4 Training and Evaluation Protocol
  • Absolute Trajectory Error (ATE): The root-mean-square error between the ground-truth and predicted 3D positions over the entire trajectory, measured in meters.
  • ATE%: The ATE normalized by the total trajectory length and expressed as a percentage ($100\times\text{ATE}/\text{length}$). This metric provides a scale-invariant measure of error, which is crucial for fairly comparing performance across our aggressive flight trajectories of varying lengths.
  • Relative Pose Error (RPE): The error measured over fixed-length sub-trajectories, capturing local drift.
@@
G.5 Eigenvector Canonicalization for the VN Baseline
To resolve ambiguities in the eigendecomposition $C=V\Lambda V^{\top}$ for the VN baseline, we canonicalize the eigenvector matrix $V=[\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}]$ as follows:
@@ G.5 Eigenvector Canonicalization for the VN Baseline
    Right-handed Frame: If $\det V<0$, we set $\mathbf{e}_{3}\leftarrow-\mathbf{e}_{3}$ to ensure $\det V=+1$.
@@
    Sign Disambiguation: For each eigenvector $\mathbf{e}_{i}$ with a distinct eigenvalue, we enforce a consistent sign by ensuring $\mathbf{v}^{\top}\mathbf{e}_{i}\geq 0$. If not, we set $\mathbf{e}_{i}\leftarrow-\mathbf{e}_{i}$.
@@
    Multiplicity Handling: In the rare case of repeated eigenvalues, we use the projection of the velocity vector $\mathbf{v}$ onto the corresponding eigenspace to deterministically define the first basis vector, then complete the basis via Gram-Schmidt.
All steps use only equivariant operations, preserving the overall symmetry of the baseline.
Appendix H Proof of $\mathrm{SO}(3)$-Equivariance for ReLN Velocity Extract with Covariance Inputs
@@
H.1 $\mathrm{SO}(3)$-Equivariant Vector Extraction via Skew-Symmetric Projection
Our network, $\Phi$, is designed to be adjoint-equivariant. It maps geometric inputs—such as a embedded velocity $K\in\mathfrak{so}(3)$ and a covariance matrix $S\in\mathrm{SPD}(3)$—to a matrix feature $A\in\mathbb{R}^{3\times 3}$. The inputs transform under the adjoint action of any rotation $R\in\mathrm{SO}(3)$:
@@ H.1 $\mathrm{SO}(3)$-Equivariant Vector Extraction via Skew-Symmetric Projection
Proposition H.1 (Equivariance of Skew-Symmetric Extraction).
Let a network $\Phi$ and its inputs transform according to Eqs. 47 and 48. If a vector $\hat{\mathbf{v}}\in\mathbb{R}^{3}$ is extracted from the output matrix $A=\Phi(K,S)$ via the projection
$A_{\mathrm{skew}}=\tfrac{1}{2}(A-A^{\top}),\qquad\hat{\mathbf{v}}=(A_{\mathrm{skew}})^{\vee},$ (49)
then the vector $\hat{\mathbf{v}}^{\prime}$ extracted from the transformed output $A^{\prime}=\Phi(K^{\prime},S^{\prime})$ transforms covariantly as $\hat{\mathbf{v}}^{\prime}=R\hat{\mathbf{v}}$.
@@
By the adjoint-equivariance property in Eq. 48, the network satisfies $\Phi(RKR^{\top},RCR^{\top})=R\,\Phi(K,C)\,R^{\top}=RAR^{\top}$. Let $A^{\prime}=RAR^{\top}$. The skew-symmetric component of the transformed output $A^{\prime}$ is:
@@
$\hat{\mathbf{v}}^{\prime}=(A^{\prime}_{\mathrm{skew}})^{\vee}=(\mathrm{Ad}_{R}(A_{\mathrm{skew}}))^{\vee}=R\,(A_{\mathrm{skew}})^{\vee}=R\hat{\mathbf{v}}$. (51)
Remark 4.
@@
Large Language Model (LLM) Usage
We used a large language model (LLM) to aid in polishing the writing and improving grammatical clarity of the manuscript. The LLM did not contribute to the research ideation, experiments, or technical content; all scientific claims and results were generated solely by the authors.

### v2

@@ H.2 Group Action Equivariance via Embedding Map
Theorem H.2 (Adjoint Equivariance).
For any $p\in\mathbb{R}^{4}$ and $\Lambda\in\mathrm{SO}^{+}(1,3)$, let $G=\operatorname{diag}(\Lambda,1)\in\operatorname{GL}(5)$. Then
@@
Remark H.3 (Generalization to Orthogonal Groups).
This embedding technique is not limited to the Lorentz group and can be readily generalized to any orthogonal group $\operatorname{O}(n)$ or special orthogonal group $\operatorname{SO}(n)$. For instance, in applications involving 3D point clouds where the symmetry is $\operatorname{SO}(3)$, a vector $p\in\mathbb{R}^{3}$ would be embedded into the Lie algebra $\mathfrak{gl}(4)$ as:
@@
H.3 Experimental Implementation
Dataset.
@@ H.3 Experimental Implementation
Model.
[…]
$m_{ij}=\phi_{e}\Big(h_{i},h_{j},\psi\big(\widetilde{B}(\varphi(p_{i}),\varphi(p_{i}))\big),\psi\big(\widetilde{B}(\varphi(p_{j}),\varphi(p_{j}))\big),\psi\big(\widetilde{B}(\varphi(p_{i}),\varphi(p_{j}))\big)\Big)$ (50)
@@
Training Setup.
For a fair comparison, our training procedure closely follows the protocol established in the LorentzNet (Gong et al., 2022). The model was trained for a total of 35 epochs on a NVIDIA RTX 4090 GPU. We used the AdamW optimizer with a weight decay of 0.01 and a batch size of 128, matching the total effective batch size from the reference work. The learning rate was managed by the paper’s specific three-stage schedule: a 4-epoch linear warm-up to an initial rate of $1\times 10^{-3}$, followed by a 28-epoch CosineAnnealingWarmRestarts schedule, and a final 3-epoch exponential decay. After each epoch, the model with the highest validation accuracy was saved for final evaluation on the test set.
Appendix I Drone Experiment Details
@@
I.1 Geometric Framework for Equivariant Covariance Processing
@@
Theorem I.1 (Equivariance of the Logarithmic Map).
@@
I.3 Baseline and Model Implementation Details
@@
Eigendecomposition-based $\mathbf{SO}\textbf{(3)}$-Equivariant Baseline.
@@
Spherical Harmonic (SH)-based Equivariant Baselines (TFN and SE(3)-Transformer).
We evaluate two representative $\mathrm{SO}(3)$-equivariant graph architectures built from irrep-valued features and steerable bases expanded in spherical harmonics (SH): Tensor Field Networks (TFN) (Thomas et al., 2018) and the SE(3)-Transformer (Fuchs et al., 2020). Both enforce equivariance by restricting linear maps to $\mathrm{SO}(3)$ intertwiners, implemented through SH steerable kernels and tensor-product (Clebsch–Gordan) couplings.
Transformation Rules and Irrep Decomposition. Under a rotation $R\in\mathrm{SO}(3)$, the velocity transforms as $\mathbf{v}\mapsto R\mathbf{v}$, while a covariance transforms by congruence $C\mapsto RCR^{\top}$. We represent the symmetric rank-2 covariance via the standard irrep decomposition $\mathrm{Sym}^{2}(\mathbb{R}^{3})\cong L{=}0\oplus L{=}2$, i.e., $C=\tfrac{\mathrm{tr}(C)}{3}I+C_{0}$ with $\mathrm{tr}(C)$ as a type-0 scalar and $C_{0}:=C-\tfrac{\mathrm{tr}(C)}{3}I$ as a type-$2$ traceless-symmetric tensor.
Feature construction and temporal graph. Each timestamp is encoded as an $\mathrm{SO}(3)$ fiber with degrees $L\in\{0,1,2\}$: $\mathrm{tr}(C)$ for L=0L{=}0, $\mathbf{v}$ for L=1L{=}1, and $C_{0}$ for L=2L{=}2. We form a temporal chain graph connecting $t$ to $t\pm 1$ and drive message passing using edge attributes derived from relative offsets. TFN aggregates messages via steerable kernel filtering (with learned radial profiles), whereas SE(3)-Transformer replaces fixed kernel aggregation with multi-head, content-dependent equivariant self-attention (Thomas et al., 2018; Fuchs et al., 2020). In both cases, uncertainty is stored in separate type-0 and type-$2$ blocks, and interactions between them are mediated by learned equivariant tensor-product couplings rather than by treating $C$ as a single matrix feature throughout the network.
Relation to our formulation. In contrast, ReLN embeds $(\mathbf{v},C)$ into a unified matrix space $\mathfrak{gl}(3)$ and processes both under a single adjoint action, directly matching the congruence rule $C\mapsto RCR^{\top}$ in a shared Lie-algebraic backbone.
Empirical observation. In our experiments, TFN consistently outperformed the SE(3)-Transformer. We attribute this to the nature of drone state estimation as a path integration task. The SH-based convolutional kernels in TFN act as learned local integrators, which are better suited for capturing high-frequency temporal dynamics than the global dependency focus of attention mechanisms. While attention provides flexibility, it can introduce susceptibility to high-frequency noise in aggressive flight sequences, whereas the local inductive bias of SH-convolutions provides a more stable prior for trajectory reconstruction.
@@
The ReLN model shares a similar backbone but incorporates the ReLN-Bracket layer. In contrast to the VN and graph-based baselines, ReLNs provide a unified framework for velocity and covariance processing within the Lie algebra $\mathfrak{gl}(3)$. Velocities $\mathbf{v}\in\mathbb{R}^{3}$ are lifted into $\mathfrak{so}(3)\subset\mathfrak{gl}(3)$ via $K=\mathbf{v}^{\wedge}$, while covariance $C$ (or $\log C$) is treated as a structured geometric input. Both transform under the same adjoint action: $K^{\prime}=RKR^{\top}$ and $C^{\prime}=RCR^{\top}$. By utilizing the matrix logarithm, $\log C\in\mathrm{Sym}(3)\subset\mathfrak{gl}(3)$, we embed the manifold-valued uncertainty into a linear subspace compatible with Lie-algebraic processing. The final velocity estimate is equivariantly extracted via the projection $\tilde{\mathbf{v}}=(\tfrac{1}{2}(A-A^{\top}))^{\vee}$ from the network’s matrix output $A\in\mathbb{R}^{3\times 3}$.
I.4 Training and Evaluation Protocol
Problem Formulation.
@@ I.4 Training and Evaluation Protocol
  • Absolute Trajectory Error (ATE): The root-mean-square error (RMSE) between the ground-truth positions $p_{k}$ and predicted 3D positions $\hat{p}_{k}$ over the entire trajectory, calculated as $\sqrt{\frac{1}{N}\sum_{k=1}^{N}\|\hat{p}_{k}-p_{k}\|^{2}}$, measured in meters.
  • ATE%: The ATE normalized by the total trajectory length $L$, expressed as a percentage ($100\times\text{ATE}/L$). This metric provides a scale-invariant measure of error, which is crucial for fairly comparing performance across our aggressive flight trajectories of varying lengths.
  • Relative Translation Error (RTE): This metric evaluates local consistency by measuring the relative translational error over fixed-length sub-trajectories. For each window of duration $\Delta t=2.0\,\text{s}$ ($n=400$ samples), we compute the RMSE of the positional difference between estimated and ground-truth relative displacements: $\|(\hat{p}_{k+n}-\hat{p}_{k})-(p_{k+n}-p_{k})\|$. We report the mean error by sliding this window across the trajectory with a stride of $0.5\,\text{s}$ ($100$ samples) to ensure a robust assessment of short-term stability.
@@
I.5 Eigenvector Canonicalization for the VN Baseline
To resolve ambiguities in the eigendecomposition $C=V\Lambda V^{\top}$ for the VN baseline while preserving $SO(3)$ geometry, we canonicalize the eigenvector matrix $V=[\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}]$ using the following sequence. To avoid producing reflection matrices (where $\det V=-1$), sign disambiguation must precede the enforcement of a right-handed frame.
@@ I.5 Eigenvector Canonicalization for the VN Baseline
    Sign Disambiguation: For each eigenvector $\mathbf{e}_{i}$ associated with a distinct eigenvalue, we enforce a consistent orientation relative to the velocity vector $\mathbf{v}$ by ensuring $\mathbf{v}^{\top}\mathbf{e}_{i}\geq 0$. If $\mathbf{v}^{\top}\mathbf{e}_{i}<0$, we set $\mathbf{e}_{i}\leftarrow-\mathbf{e}_{i}$.
@@
    Right-handed Frame: After the individual signs are fixed, we ensure the matrix $V$ represents a proper rotation. If $\det V<0$, we flip the sign of the third eigenvector, $\mathbf{e}_{3}\leftarrow-\mathbf{e}_{3}$, to ensure $\det V=+1$. This step is performed last to ensure that any sign changes in the previous step do not inadvertently result in a reflection frame.
@@
    Multiplicity Handling: In cases of repeated eigenvalues, we project the velocity vector $\mathbf{v}$ onto the corresponding eigenspace to uniquely define the first basis vector of that subspace, then complete the orthonormal basis via the Gram-Schmidt process, followed by the determinant check in Step 2.
By applying the determinant constraint after sign alignment, we guarantee that the baseline operates on valid rotation matrices, ensuring a fair and geometrically consistent comparison with our $SO(3)$-equivariant model.
I.6 Full Experimental Results and Ablation Study for Drone State Estimation
Table 13 provides the exhaustive results for all models across three input modalities: $v,(v,C),(v,\log C)$. This comprehensive comparison highlights that while the $\log$-covariance interface generally improves performance for most equivariant models, our ReLN architecture remains the superior backbone due to its ability to jointly model the full reductive structure.
Table 13: Complete performance report on the drone dataset including all input variants and ablations. We report ATE (meters), ATE% (scaled), and RTE (meters). SO(3) denotes evaluation under random test-time rotations.
Model Input ID (In-Distribution) SO(3) (Rotated)
ATE $\downarrow$ ATE% $\downarrow$ RTE $\downarrow$ ATE $\downarrow$ ATE% $\downarrow$ RTE $\downarrow$
ResNet $v$ 208.07 95.06 107.60 217.02 100.39 111.29
ResNet $(v,C)$ 205.11 94.94 106.07 213.26 98.90 109.37
Vector Neurons $v$ 17.36 7.52 13.51 17.36 7.52 13.51
Vector Neurons $(v,C)$ 191.78 88.66 98.39 190.22 88.47 98.26
Tensor Field Network $v$ 24.59 10.95 18.23 24.59 10.95 18.23
Tensor Field Network $(v,C)$ 17.56 7.60 14.40 17.56 7.60 14.40
Tensor Field Network $(v,\log C)$ 16.83 7.56 13.34 16.83 7.56 13.34
SE(3)-Transformer $v$ 22.22 9.85 17.63 22.22 9.85 17.63
SE(3)-Transformer $(v,C)$ 21.67 9.37 16.77 21.67 9.37 16.77
SE(3)-Transformer $(v,\log C)$ 20.12 8.84 15.36 20.12 8.84 15.36
Ablations (Ours)
Lie Neurons $(v,C)$ 16.86 7.43 13.65 16.86 7.43 13.65
Lie Neurons $(v,\log C)$ 15.65 6.76 12.04 15.65 6.76 12.04
ReLN (no semisimple) $(v,\log C)$ 16.27 7.00 12.65 16.27 7.00 12.65
Our Equivariant Model
ReLN $v$ 16.85 7.31 12.70 16.85 7.31 12.70
ReLN $(v,C)$ 16.49 7.21 13.02 16.49 7.21 13.02
ReLN $(v,\log C)$ 13.92 5.99 11.04 13.92 5.99 11.04
The Lie Neurons architecture (Lin et al., 2024a) is mathematically equivalent to the “no-center” variant of ReLN, i.e., ReLN with the bilinear form restricted to the semisimple ideal.
Figure 7: Qualitative comparison of reconstructed trajectories ReLN variants maintain superior tracking fidelity and stability compared to non-equivariant and standard equivariant baselines.
Qualitative Analysis of Trajectory Fidelity.
Figure 7 validates the qualitative robustness of ReLNs under high-dynamic flight conditions. While standard MLP-based ResNets and eigendecomposition-based Vector Neurons suffer from rapid error accumulation and structural divergence, ReLN maintains high fidelity to the ground truth even in velocity-only configurations.
The integration of covariance, $(v,C)$, and specifically the $\log$-covariance provides effective regularization via uncertainty-aware weighting. Notably, while steerable baselines such as TFN and SE(3)-Transformers show improved stability with $\log$-covariance, they still exhibit observable deviations during sharp maneuvers where geometric consistency is critical. This confirms our finding that the reductive Lie-algebraic backbone offers a more stable and geometrically principled space for fusing heterogeneous sensor signals and their associated uncertainties.
Appendix J Proof of $\mathrm{SO}(3)$-Equivariance for ReLN Velocity Extract with Covariance Inputs
@@ I.6 Full Experimental Results and Ablation Study for Drone State Estimation
J.1 $\mathrm{SO}(3)$-Equivariant Vector Extraction via Skew-Symmetric Projection
Our network, $\Phi$, is designed to be adjoint-equivariant. It maps geometric inputs—such as an embedded velocity $K\in\mathfrak{so}(3)$ and a covariance matrix $S\in\mathrm{SPD}(3)$—to a matrix feature $A\in\mathbb{R}^{3\times 3}$. The inputs transform under the adjoint action of any rotation $R\in\mathrm{SO}(3)$:
@@
Proposition J.1 (Equivariance of Skew-Symmetric Extraction).
Let a network $\Phi$ and its inputs transform according to Eqs. 58 and 59. If a vector $\tilde{\mathbf{v}}\in\mathbb{R}^{3}$ is extracted from the output matrix $A=\Phi(K,S)$ via the projection
$A_{\mathrm{skew}}=\tfrac{1}{2}(A-A^{\top}),\qquad\tilde{\mathbf{v}}=(A_{\mathrm{skew}})^{\vee},$ (60)
then the vector $\tilde{\mathbf{v}}^{\prime}$ extracted from the transformed output $A^{\prime}=\Phi(K^{\prime},S^{\prime})$ transforms covariantly as $\tilde{\mathbf{v}}^{\prime}=R\tilde{\mathbf{v}}$.
@@
By the adjoint-equivariance property in Eq. 59, the network satisfies $\Phi(RKR^{\top},RSR^{\top})=R\,\Phi(K,S)\,R^{\top}=RAR^{\top}$. Let $A^{\prime}=RAR^{\top}$. The skew-symmetric component of the transformed output $A^{\prime}$ is:
@@
$\tilde{\mathbf{v}}^{\prime}=(A^{\prime}_{\mathrm{skew}})^{\vee}=(\mathrm{Ad}_{R}(A_{\mathrm{skew}}))^{\vee}=R\,(A_{\mathrm{skew}})^{\vee}=R\tilde{\mathbf{v}}$. (62)
Remark J.2.
@@
Appendix K Details on 3D Gaussian Splatting Experiments
In this section, we detail the experimental protocols and architectural modifications for the 3D Gaussian Splatting (3DGS) experiments. Figure 5 provides a comprehensive overview of the ReLN-integrated Gaussian-MAE architecture, highlighting the equivariant data flow from geometric lifting to projection. We provide a rigorous theoretical justification for our architectural design to achieve end-to-end equivariance by jointly learning geometric and photometric features within a unified Lie-algebraic framework.
K.1 Datasets and Protocol
We adopt a standard two-stage protocol common in self-supervised learning for 3D representation (Ma et al., 2025), separating the pre-training domain from the downstream evaluation task to assess generalization.
Pre-training: ShapeSplat (ShapeNet-derived).
We utilize ShapeSplat (Ma et al., 2025), a large-scale collection of object-level 3D Gaussian primitives derived from ShapeNet. Each object is represented by a set of anisotropic Gaussians $\Theta_{i}=\{\mu_{i},\Sigma_{i},\alpha_{i},c_{i}\}_{i=1}^{N}$. Following the baseline protocol, we sample $N=1,024$ Gaussians via furthest point sampling (FPS) and apply a $60\%$ masking ratio for the masked autoencoding objective.
Downstream: ModelNet10 Classification.
To evaluate rotational robustness, we fine-tune the pre-trained encoder on ModelNet10. We report accuracy on: (i) Aligned: The canonical test set. (ii) Rotated: The test set under random rotations $R\sim SO(3)$, testing whether the learned representations remain stable under arbitrary orientations.
K.2 Holistic Equivariance via Joint Feature Learning
A critical limitation of the baseline architecture (Ma et al., 2025) is the naive concatenation of heterogeneous Gaussian attributes. A 3D Gaussian combines equivariant quantities ($\mu,\Sigma$) with invariant quantities ($c,\alpha$). Our goal in ReLN is to construct a holistic equivariant architecture where these distinct attribute types are learned jointly without violating the symmetry of the feature space.
1. The Equivariant Component (Active Geometry).
The mean position $\mu\in\mathbb{R}^{3}$ and covariance $\Sigma\in\mathbb{R}^{3\times 3}$ dictate the spatial structure. Under a global rotation $R\in SO(3)$, these transform via $\mu\mapsto R\mu$ and $\Sigma\mapsto R\Sigma R^{\top}$. Recognizing that the covariance update corresponds to the adjoint action $\mathrm{Ad}_{R}(\Sigma)=R\Sigma R^{-1}$, we embed $\Sigma$ into the symmetric subspace of $\mathfrak{gl}(3)$. This ensures that the primary geometric features are processed within a consistent $\mathrm{Ad}$-equivariant vector space throughout the network.
2. The Invariant Component (Photometry and Geometric Norms).
To support robust representation learning, we augment the state space with a set of $O(3)$-invariant scalars (Type-0 features). This set includes:
[…]
  • Derived Geometric Invariants: Crucially, we also explicitly compute rotation-invariant geometric descriptors, such as the norms of position vectors ($\|\mu\|$) and rotation/scale magnitudes derived from $\Sigma$.
In the ReLN framework, treating these as Type-0 features serves to facilitate its feature integration. These scalars participate in the learning process by modulating the equivariant features without violating the underlying group structure and equivariance.
3. Joint Learning in ReLN Attention.
The explicit stratification of these streams is a structural prerequisite for joint equivariant learning. By assigning each attribute to its correct algebraic type (Type-1/Type-2 for geometry, Type-0 for scalars), the ReLN attention mechanism can validly integrate information across all modalities. The invariant scalars ($c,\alpha,\|\mu\|$) contribute to the computation of attention scores (invariant inner products), thereby modulating how the equivariant geometric features ($\mu,\Sigma$) are aggregated. This design ensures that the entire network optimizes a single, cohesive objective function while guaranteeing that the final representation remains equivariant end-to-end.
K.3 Architectural Modifications
To enforce this holistic equivariance, we replaced non-robust components of the baseline architecture with ReLN equivalents that respect the defined type system.
ReLNEncoder and Geometric Aggregation.
Our ReLNEncoder begins with a Geometric Lifting stage that maps raw Gaussian parameters ($C,S,R$) into $\mathfrak{gl}(3)$. The baseline uses a ‘SoftEncoder‘ (MLP on raw coordinates) which breaks equivariance because standard linear weights $W$ do not rotate with the input frame. We replaced this with ReLNEncoder, which employs $\mathrm{Ad}$-equivariant linear layers (ReLNLinear) to process local neighborhoods. ReLNEncoder incorporates a Geometric Aggregator, which utilizes an equivariant norm-based max pooling mechanism to compress grouped splats into structured $C^{\prime}\times 9$ tokens. This ensures that the fundamental geometric tokens are processed within a mathematically consistent vector space.
Central Positional Embedding.
To provide spatial context without violating symmetry, we process group centroids ($\mu$) through a dedicated Central Positional Embedding. This module lifts 3D spatial coordinates into a structured Lie-algebraic representation, which is then injected into both the encoder and decoder stages (as illustrated in Figure 5). This injection strategy ensures that the transformer blocks distinguish relative spatial relations while maintaining rotational equivariance across the entire processing pipeline.
Invariant Global Pooling (vs. CLS Token).
For downstream classification tasks, we eschew the conventional learnable [CLS] token, recognizing that a fixed-reference-frame parameter inherently imposes a canonical orientation bias that violates rotational symmetry. Instead, we employ a permutation-invariant Global Max Pooling over the final sequence of equivariant tokens. This ensures that the aggregated global representation remains stable and rotates consistently with the input geometry, thereby satisfying holistic rotational symmetry for robust 3D representation learning.
K.4 Training and Evaluation Protocol
Pre-training Setup. We pre-trained both the baseline and ReLN models for 300 epochs on the ShapeSplat (ShapeNet) dataset. We utilized the AdamW optimizer with an initial learning rate of $1\times 10^{-3}$, a weight decay of $0.05$, and a cosine annealing scheduler with a 10-epoch warm-up. For the ReLN model, we explicitly disabled the soft_knn option to prevent the leakage of non-equivariant features from the legacy encoder implementation.
Fine-tuning Setup. For the downstream ModelNet10 classification task, we fine-tuned the pre-trained encoder for an additional 300 epochs. In this phase, we adjusted the learning rate to $5\times 10^{-4}$ while maintaining the same weight decay ($0.05$) and cosine scheduling strategy.
Evaluation Metrics. To rigorously evaluate geometric stability, we construct two test benchmarks:
  • Standard (Aligned): The original ModelNet10 test set where objects are canonically aligned.
  • Rotated (SO(3)): We apply a random rotation $R\in\mathrm{SO}(3)$ to every object. For a Gaussian splat, we transform parameters as $\mu^{\prime}=R\mu$, $\Sigma^{\prime}=R\Sigma R^{\top}$.
K.5 Additional Convergence Results
Figure 8 presents the detailed training dynamics. The ReLN model consistently achieves lower loss values across all geometric attributes—Rotation, Scale, and Spherical Harmonics (SH)—demonstrating that enforcing equivariance stabilizes the learning of intrinsic 3D shape descriptors.
Figure 8: Pre-training Dynamics on 3D Gaussian Splats. Comparison of training convergence between the Baseline (Ma et al., 2025) and ReLN (Ours) across 300 epochs on ShapeNet. The red shaded regions highlight the gap where ReLN achieves lower reconstruction error than the baseline. ReLN consistently outperforms the baseline across all tracked metrics (including Rotation, Scale, Density, and Chamfer Distance), demonstrating that equivariance improves every aspect of 3D Gaussian reconstruction.
Appendix L Experimental Details for the Double-Pendulum Benchmark
To ensure a rigorous and fair comparison, we adhere to the experimental protocol and system parameters of the EMLP benchmark (Finzi et al., 2021).
System dynamics and dataset generation.
We learn Hamiltonian dynamics of a 3D double spring pendulum exhibiting $O(2)$ symmetry about the $z$-axis. All physical constants are set to unity ($m_{i}=k_{i}=l_{i}=1.0$) with gravity $g=[0,0,1]$. Ground-truth trajectories are generated by Hamiltonian integration over a horizon of $T=30$s. We construct $1500$ trajectory chunks of duration $1$s (sampled at five $0.2$s intervals) and split them into $500/500/500$ for training/validation/testing.
Models and training.
[…]
EMLP-HNN baseline (equivariant layer construction).
Let $G=O(2)$ be the symmetry group acting on the $(x,y)$ plane (with $z$ unchanged), e.g., $R(\phi)=\mathrm{diag}(R_{2\times 2}(\phi),1)$ (and similarly for reflections), and let $\rho(\cdot)$ denote the induced representation on the state space. Equivariance of a linear map $W:V_{\mathrm{in}}\!\to\!V_{\mathrm{out}}$ requires
$\rho_{\mathrm{out}}(g)\,W\;=\;W\,\rho_{\mathrm{in}}(g),\qquad\forall g\in G.$ (63)
Following (Finzi et al., 2021), EMLP enforces this by constraining the vectorized weights $v=\mathrm{vec}(W)$ under the product representation $\rho=\rho_{\mathrm{out}}\otimes\rho_{\mathrm{in}}^{*}$:
$\rho(g)\,v\;=\;v,\qquad\forall g\in G.$ (64)
To avoid infinitely many constraints, the condition is reduced to a finite generator set consisting of Lie algebra generators $\{A_{i}\}$ and discrete generators $\{h_{j}\}$:
$d\rho(A_{i})\,v=0\quad\text{and}\quad(\rho(h_{j})-I)\,v=0,$ (65)
which are assembled into a linear system $Cv=0$. An equivariant basis $Q$ is obtained by computing a nullspace of $C$ (e.g., via SVD/Krylov methods as in the reference implementation), and the layer is parameterized as $v=Q\beta$ during training (no further constraint solving during SGD). This basis computation is group- and representation-specific (e.g., $O(2)$ vs. $SO(2)$).
EMLP nonlinearity.
We use the standard EMLP gated nonlinearity (Finzi et al., 2021), where scalar channels produce gates that multiplicatively modulate non-scalar feature blocks, preserving equivariance while enabling expressive nonlinear mixing.
ReLN-HNN (ours).
In contrast, ReLN-HNN bypasses numerical constraint solving and constructs equivariant/invariant primitives in closed form within a reductive Lie-algebraic space (here $\mathfrak{gl}(3)$), using bilinear/Killing-form-based operations to obtain symmetry-consistent features without group-specific basis recomputation. The same architecture is used across different symmetry settings (e.g., $O(2)$, $SO(2)$, $D_{6}$) without re-solving linear constraints.
Computational efficiency measurement.
Per-step FLOPs in Table 6 are measured using JAX HLO cost analysis, which counts device-executed floating-point operations including internal tensor contractions and projections.